find the x intercept, if any, and y intercept, if any, of the line, use the intercepts to sketch a graph of the line and find the slope of the line.
2x + 7y = 0

Answer by ewatrrr(24785)

 
Hi,
find the x intercept, if any, and y intercept, if any, of the line, use the intercepts
to sketch a graph of the line and find the slope of the line.
2x + 7y = 0 | y-intercept is (0,0) x-intercept is (0,0)
Using the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept.
2x + 7y = 0 OR y = (-2/7)x
Note: as y decreases by 2, x increases by 7. Ex Pt(7,-2) on this Line